OpenStudy (anonymous):

Black: 3 Gray stripes: 5 White: 2 A. If Sally chooses 1 kitten at random to place in the window, what is the probability that she will choose a black kitten? Explain your answer. B. If Sally chooses 2 kittens at random for the window, what is the probability that the first kitten will be black and the second kitten will be white? Explain your answer.

OpenStudy (ashely14):

3 out of 7

OpenStudy (cecil_the_weasel):

A.) There are 10 kittens total and 3 are black. So 3/10

OpenStudy (ashely14):

oh yeah lol i forgot to add the 3

OpenStudy (anonymous):

I need help on B

OpenStudy (cecil_the_weasel):

B.) The probability of choosing a black kitten is 3/10. The probability of choosing a white kitten the second time is 2/9 because there are 9 left to choose from. Can you figure out what is next?

OpenStudy (cecil_the_weasel):

2/9 is not the answer

OpenStudy (cecil_the_weasel):

You still there Batman?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Can you still help me more with B @cecil_the_weasel

OpenStudy (cecil_the_weasel):

Sure ☺

OpenStudy (cecil_the_weasel):

So the probability of the first event (choosing a black cat) is 3/10. Lets call this 'A'.The probability of the second event (picking a white cat) is 2/9. Lets call this 'B'. Now to find P(A then B) you multiply the two probabilities together.

OpenStudy (cecil_the_weasel):

\[\frac{ 3 }{10}\times \frac{ 2 }{ 9 }\]

OpenStudy (anonymous):

So I get 6/90 @cecil_the_weasel

OpenStudy (cecil_the_weasel):

Yes. Which can be reduced to?

OpenStudy (anonymous):

1/15 @cecil_the_weasel

OpenStudy (cecil_the_weasel):

Great job! There's your answer :)

OpenStudy (anonymous):

Thanks @cecil_the_weasel

OpenStudy (anonymous):

How do I explain B tho @cecil_the_weasel

OpenStudy (cecil_the_weasel):

Sorry, I had to take a test so it took me a while to get back. I am never very good at answering that but I found if you just tell how you solved it you will likely get full credit ☺